mw104969:

A single batch of cleaning solution contains 3/4 cup of bleach. How many cups of bleach are in 5 batches of the cleaning solution? 3/20 15/20 1 1/4 2 3 3/4

swinn:

all you need to do is multiply 3/4*5

mw104969:

i did but i didnt get the right answer

swinn:

do you have options?

mw104969:

the options are above but when i multiplied the equation i got 15/4

swinn:

then all you need to do now is convert the improper fraction into a fixed number

mw104969:

so it would be \[3 3/4\]

mw104969:

3 3/4*

swinn:

3/4*5/11=15/4= 3 3/4 Result in decimals: 3.75 Calculation steps: 3/4×5/1= 3 × 5 4 × 1= 15/4=3 3/4

swinn:

so yes you are correct

mw104969:

thank you

swinn:

np